# Optimization of foreground moment deprojection for semi-blind CMB polarization reconstruction

### v1

@@ 4.2 Optimization of moment deprojection
Table 2: Optimal configuration of moments and dust pivot parameters for deprojection, as determined by ocMILC across various sky regions and needlet scales for each foreground data set. The optimal deprojection coefficient for each moment SED is indicated in parentheses. As detailed in Section 4.2, the optimal combination of these parameters is chosen by minimizing the denoised output variance of Equation 4.12.
Band $m_{\textrm{fgds}}$ d1s1 d4s1 d7s1 d12s1
$j=1$ $5$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=19.0$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.001)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{s}}}(-0.005)\big{]}$
$6$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}f_{\textrm{d}}(0.001)$,$f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{s}}}(0.)$, $\partial_{T_{\textrm{d}}}(0.)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.01)$,$\partial^{2}_{T_{\textrm{d}}}(0.01)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.25,\ \bar{T}_{\textrm{d}}=18.25$ $\big{[}f_{\textrm{d}}(0.)$,$f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{d}}}(-0.005)$, $\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$
$j=2$ $4$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.5$; $\big{[}f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{s}}}(-0.0005)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.0005)$,$\partial_{\beta_{\textrm{s}}}(0.0005)$, $\partial_{T_{\textrm{d}}}(-0.001)$,$\partial^{2}_{\beta_{\textrm{d}}}(0.01)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(-0.001)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$
$5$ $\bar{\beta}_{\textrm{d}}=1.7,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{d}}(0.)$,$\partial_{\beta_{\textrm{d}}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial^{2}_{\beta_{\textrm{d}}}(0.01)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=19.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.005)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=21.5$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.001)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.7,\ \bar{T}_{\textrm{d}}=17.5$ $\big{[}f_{\textrm{s}}(-0.0005)$,$\partial_{\beta_{\textrm{s}}}(0.001)$, $\partial_{T_{\textrm{d}}}(0.),$ $\partial^{2}_{\beta_{\textrm{d}}}(-0.0005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$
$j=3$ $3$ $\bar{\beta}_{\textrm{d}}=1.5,\ \bar{T}_{\textrm{d}}=19.0$; $\big{[}f_{\textrm{s}}(-0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(-0.001)$, $\partial_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=21.0$; $\big{[}f_{\textrm{s}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$
$4$ $\bar{\beta}_{\textrm{d}}=1.65,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}\partial_{\beta_{\textrm{d}}}(0.001)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=20.0$; $\big{[}\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=19.25$ $\big{[}\partial_{\beta_{\textrm{s}}}(0.005)$,$\partial^{2}_{\beta_{\textrm{d}}}(-0.0005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$
$5$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=19.0$ $\big{[}\partial_{\beta_{\textrm{d}}}(-0.001)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$
$j=4$ $2$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(-0.005)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=18.25$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.65,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}f_{\textrm{d}}(-0.01)$,$f_{\textrm{s}}(-0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(-0.001)\big{]}$
$3$ $\bar{\beta}_{\textrm{d}}=1.25,\ \bar{T}_{\textrm{d}}=17.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=17.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=20.25$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.5,\ \bar{T}_{\textrm{d}}=21.25$ $\big{[}f_{\textrm{s}}(-0.001)$,$\partial_{\beta_{\textrm{d}}}(-0.0005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$
$4$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=17.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(0.0005)$,$\partial_{\beta_{\textrm{s}}}(0.001)$, $\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=20.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.001)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(-0.001)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=19.5$; $\big{[}\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=17.0$ $\big{[}\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$
$j=5$ $1$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}\partial^{2}_{\beta_{\textrm{s}}}(0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(-0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$
$2$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(0.0005),\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{d}}(0.001),f_{\textrm{s}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(0.),\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(0.001),\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(-0.001)\big{]}$
$3$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{s}}(0.001)$,$\partial_{\beta_{\textrm{d}}}(0.001)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{s}}}(0.001)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=21.5$ $\big{[}f_{\textrm{s}}(0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$
$4$ $\bar{\beta}_{\textrm{d}}=1.55,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{s}}(-0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.001)$, $\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}f_{\textrm{s}}(0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.0005)$, $\partial_{T_{\textrm{d}}}(0.0005)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{s}}(-0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.005)$, $\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=17.0$ $\big{[}\partial_{\beta_{\textrm{s}}}(0.01)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$

### v2

@@ 4.2.2 CMB $B$-mode reconstruction
Table 2: Optimal configuration of moments and dust pivot parameters for deprojection, as determined by ocMILC across various sky regions and needlet scales for each foreground data set. The optimal deprojection coefficient for each moment SED is indicated in parentheses. As detailed in Section 3.2.2, the optimal combination of these parameters is chosen by minimizing the denoised output variance of Equation 3.24.
{adjustwidth}
-1.5cm             Band $m_{\textrm{fgds}}$ d1s1 d4s1 d7s1 d12s1 $j=1$ $5$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=19.0$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.001)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{s}}}(-0.005)\big{]}$ $6$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}f_{\textrm{d}}(0.001)$,$f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{s}}}(0.)$, $\partial_{T_{\textrm{d}}}(0.)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.01)$,$\partial^{2}_{T_{\textrm{d}}}(0.01)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.25,\ \bar{T}_{\textrm{d}}=18.25$ $\big{[}f_{\textrm{d}}(0.)$,$f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{d}}}(-0.005)$, $\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$ $j=2$ $4$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.5$; $\big{[}f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{s}}}(-0.0005)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.0005)$,$\partial_{\beta_{\textrm{s}}}(0.0005)$, $\partial_{T_{\textrm{d}}}(-0.001)$,$\partial^{2}_{\beta_{\textrm{d}}}(0.01)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(-0.001)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $5$ $\bar{\beta}_{\textrm{d}}=1.7,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{d}}(0.)$,$\partial_{\beta_{\textrm{d}}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial^{2}_{\beta_{\textrm{d}}}(0.01)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=19.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.005)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=21.5$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{s}}}(0.001)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.7,\ \bar{T}_{\textrm{d}}=17.5$ $\big{[}f_{\textrm{s}}(-0.0005)$,$\partial_{\beta_{\textrm{s}}}(0.001)$, $\partial_{T_{\textrm{d}}}(0.),$ $\partial^{2}_{\beta_{\textrm{d}}}(-0.0005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$ $j=3$ $3$ $\bar{\beta}_{\textrm{d}}=1.5,\ \bar{T}_{\textrm{d}}=19.0$; $\big{[}f_{\textrm{s}}(-0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(-0.001)$, $\partial_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=21.0$; $\big{[}f_{\textrm{s}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $4$ $\bar{\beta}_{\textrm{d}}=1.65,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}\partial_{\beta_{\textrm{d}}}(0.001)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=20.0$; $\big{[}\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=19.25$ $\big{[}\partial_{\beta_{\textrm{s}}}(0.005)$,$\partial^{2}_{\beta_{\textrm{d}}}(-0.0005)$, $\partial^{2}_{\beta_{\textrm{s}}}(0.05)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $5$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=19.0$ $\big{[}\partial_{\beta_{\textrm{d}}}(-0.001)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $j=4$ $2$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(-0.005)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=18.25$; $\big{[}f_{\textrm{d}}(0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.65,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}f_{\textrm{d}}(-0.01)$,$f_{\textrm{s}}(-0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(-0.001)\big{]}$ $3$ $\bar{\beta}_{\textrm{d}}=1.25,\ \bar{T}_{\textrm{d}}=17.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=17.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=20.25$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.5,\ \bar{T}_{\textrm{d}}=21.25$ $\big{[}f_{\textrm{s}}(-0.001)$,$\partial_{\beta_{\textrm{d}}}(-0.0005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $4$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=17.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(0.0005)$,$\partial_{\beta_{\textrm{s}}}(0.001)$, $\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=20.5$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.001)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(-0.001)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=19.5$; $\big{[}\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.0005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=17.0$ $\big{[}\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$ $j=5$ $1$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}\partial^{2}_{\beta_{\textrm{s}}}(0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(-0.005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.45,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $2$ $\bar{\beta}_{\textrm{d}}=1.2,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(0.0005),\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{d}}(0.001),f_{\textrm{s}}(0.0005)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(0.),\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(0.001),\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(-0.001)\big{]}$ $3$ $\bar{\beta}_{\textrm{d}}=1.75,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{s}}(0.001)$,$\partial_{\beta_{\textrm{d}}}(0.001)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.4,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}\partial_{\beta_{\textrm{d}}}(-0.01)$,$\partial_{\beta_{\textrm{s}}}(0.01)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.35,\ \bar{T}_{\textrm{d}}=17.0$; $\big{[}f_{\textrm{s}}(0.)$,$\partial_{\beta_{\textrm{s}}}(0.001)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=21.5$ $\big{[}f_{\textrm{s}}(0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $4$ $\bar{\beta}_{\textrm{d}}=1.55,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{s}}(-0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.001)$, $\partial_{\beta_{\textrm{s}}}(0.001)$,$\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.3,\ \bar{T}_{\textrm{d}}=17.25$; $\big{[}f_{\textrm{s}}(0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.0005)$, $\partial_{T_{\textrm{d}}}(0.0005)$, $\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=21.75$; $\big{[}f_{\textrm{s}}(-0.0005)$,$\partial_{\beta_{\textrm{d}}}(-0.005)$, $\partial_{\beta_{\textrm{s}}}(0.005)$, $\partial^{2}_{T_{\textrm{d}}}(0.)\big{]}$ $\bar{\beta}_{\textrm{d}}=1.6,\ \bar{T}_{\textrm{d}}=17.0$ $\big{[}\partial_{\beta_{\textrm{s}}}(0.01)$,$\partial_{T_{\textrm{d}}}(0.)$, $\partial^{2}_{\beta_{\textrm{d}}}(0.01)$,$\partial^{2}_{\beta_{\textrm{d}}T_{\textrm{d}}}(0.)\big{]}$